OpenStudy (anonymous):

use the angle identity to find the exact value of each: cos 105 degrees and sin 195 degrees

OpenStudy (anonymous):

@ikram002p do you know how to solve this question

OpenStudy (anonymous):

for the first one, this is how i solved it, ' cos105=cos(60+45)=cos60*cos45+sin60*sin45 =sqrt(1)/2*sqrt(2)/2+sqrt(3)/2*sqrt(2) =sqrt(2)+sqrt(6)/4

OpenStudy (ikram002p):

105=45+60 so cos 105=cos( 45+60)=cos 60 cos 45 - sin 60 sin 45

OpenStudy (ikram002p):

yeah ur correct

OpenStudy (anonymous):

but its wrong, do we not square

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

because i'm evaluating my functions incorrectly

OpenStudy (ikram002p):

one sec there is negative sighn not positive

OpenStudy (anonymous):

why

OpenStudy (anonymous):

if something is not squared and something is, we cannot add it

OpenStudy (ikram002p):

cos (A+B)= cos A cos B - sin A sin B

OpenStudy (anonymous):

then it would be 1/2*sqrt(2)/2-sqrt(3)/2*sqrt(2)/2

OpenStudy (anonymous):

how would we add it up

OpenStudy (ikram002p):

1/4 sqrt 2 - sqrt 6 /4 = (sqrt 2- sqrt 6)/4 leave it like this

OpenStudy (anonymous):

we do not subtract it

OpenStudy (anonymous):

what about sin 195

OpenStudy (anonymous):

this is how i solved it 2) Sin195=Sin(150+45)=Sin150*Cos45+Sin150*cos45 =1/2*sqrt(2)/2+1/2*sqrt(2)/2 =sqrt(2)+sqrt(2)/4

OpenStudy (ikram002p):

nope sin (A+B)= sin Acos B + cos A sin B

OpenStudy (anonymous):

cos 75 degrees cos75=cos(45+30)=cos45*cos30-sin45*sin30 =sqrt(2)/2*sqrt(3)/2+sqrt(2)/2*1/2 =sqrt(6)+sqrt(2)/4 (EXACT VALUE)

OpenStudy (ikram002p):

thats right

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

sin105=sin(90+45)=sin90*cos45+Sin90+cos45 =1/1*sqrt(2)/2+1/1*sqrt(2)/2 =sqrt(2)+sqrt(2)/2 (EXACT VALUE)

OpenStudy (ikram002p):

nope

OpenStudy (anonymous):

where is my fault

OpenStudy (ikram002p):

sin 1505= sin 60 cos 45+cos 60 sin 45

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

) Show sin 4θ = 4sin θ cos θ cos 2 θ Sin2 θ=2sin(2 θ)cos(2 θ) (4sin xcos x)cos2 x =2(2sin xcos x)cos2 x =2sin2 xcos2 x Put 2x= θ =2sin θ cos θ =sin 2 θ , θ=2x =sin4x QED

OpenStudy (anonymous):

i solved sin 105

OpenStudy (anonymous):

@ikram002p would sin theta equation above be correct

OpenStudy (ikram002p):

i dnt understand ur question ?!

OpenStudy (anonymous):

Show sin 4θ = 4sin θ cos θ cos 2 θ

OpenStudy (anonymous):

i solved it above

OpenStudy (ikram002p):

sin 4θ = 4sin θ cos θ cos 2 θ sin 4θ =(sin 2 θ +2 θ) =sin 2 θ cos 2 θ+sin 2 θ cos 2 θ =2 sin 2 θ cos 2 θ =2(2 sin θ cos θ)cos 2 θ =4 sin θ cosθ cos 2 θ

OpenStudy (ikram002p):

this is how u show it , btw i coudnt go through ur steps